OpenStudy (anonymous):

A charity fair raised $6,000 by selling 500 lottery tickets. There were two types of lottery tickets: A tickets cost $10 each, and B tickets cost $60 each. How many tickets of each type were sold?

OpenStudy (anonymous):

"hit me baby one more time"

OpenStudy (anonymous):

hahah

OpenStudy (anonymous):

A: 250,250 b:315,185 C370,230 D:480,20

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

So the equations are 10a + 60b = 6000 and a + b = 500 You could just solve those equations, but I would guess and check on a multiple choice test. 6000/500 = 12, which is so close to 10 that a large majority of the tickets sold were 'a' tickets. 10(480) + 60(20) = 4800 + 1200 = 6000 bingo. The answer is c. 480, 20

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome!! Xo!

OpenStudy (anonymous):

can you help with this.... A school arranges for buses and vans to transport 48 students to the school picnic. A bus can accommodate 17 people, and a van can accommodate 9 people. Two teachers are supposed to travel in a bus, and 1 teacher is supposed to travel in a van. How many buses and vans will be required to accommodate all 53 people (students and teachers) with no seat left vacant?